# PlagBench: Exploring the Duality of Large Language Models in Plagiarism Generation and Detection

### v1

@@ Abstract
Recent literature has highlighted potential risks to academic integrity associated with large language models (LLMs), as they can memorize parts of training instances and reproduce them in the generated texts without proper attribution. In addition, given their capabilities in generating high-quality texts, plagiarists can exploit LLMs to generate realistic paraphrases or summaries indistinguishable from original work. In response to possible malicious use of LLMs in plagiarism, we introduce PlagBench, a comprehensive dataset consisting of 46.5K synthetic plagiarism cases generated using three instruction-tuned LLMs across three writing domains. The quality of PlagBench is ensured through fine-grained automatic evaluation for each type of plagiarism, complemented by human annotation. We then leverage our proposed dataset to evaluate the plagiarism detection performance of five modern LLMs and three specialized plagiarism checkers. Our findings reveal that GPT-3.5 tends to generates paraphrases and summaries of higher quality compared to Llama2 and GPT-4. Despite LLMs’ weak performance in summary plagiarism identification, they can surpass current commercial plagiarism detectors. Overall, our results highlight the potential of LLMs to serve as robust plagiarism detection tools.
@@ 1 Introduction
Growing popularity of large language models (LLMs) has dramatically shifted the way users craft their writings. With one button click, LLMs such as OpenAI’s ChatGPT [^1] https://chatgpt.com and Google’s Gemini Team et al. (2023) can deliver high-quality texts per users’ requests. For instance, users can ask LLMs to generate a whole story or revise a draft writing. While an impressive ability, it is important to underscore that LLMs’ increasing applications in various writing tasks is accompanied by substantial risks to privacy, intellectual property (IP), academic integrity, and social ethics. Many recent papers (e.g., Carlini et al. (2021); Tirumala et al. (2022); Zhang et al. (2023); Carlini et al. (2023)) revealed that LLMs can memorize information from training data without overfitting, including sensitive information, and reproduce it during inference. Not limited to verbatim copies, they may emit paraphrased or elongated versions of training data Lee et al. (2023b). The observed behavior is problematic because: (1) LLMs are generally trained on vast amounts of text documents without considering how the content was originally released or whether it is copyrighted Brown et al. (2022); Henderson et al. (2023), and (2) LLMs are not currently equipped with the ability to accurately credit the original source of information as humans do McGowan et al. (2023); Huang and Chang (2023). Hence, if we compare these violations to human actions, they can be viewed as plagiarismthe act of employing another person’s work or ideas without proper attribution.
@@
To address plagiarism in LLMs, it is critical to improve automatic plagiarism detectors. Current detectors are typically trained on data related to human-authored plagiarism, while machine-generated plagiarism may exhibit significant differences Becker et al. (2023). Khalil and Er (2023) show that ChatGPT outputs were prone to bypass popular plagiarism check software Khalil and Er (2023). Automatic plagiarism detection task is also essential in a context where bad actors purposely utilize LLMs to obfuscate original content with paraphrases or summaries and evade existing detectors. For instance, Krishna et al. (2024); Sadasivan et al. (2023); Weber-Wulff et al. (2023) find that a simple paraphrase of machine-written text can fool widely used AI-generated content and plagiarism detectors. These results shed light on the urgent necessity of developing more reliable, robust, and universal plagiarism detectors.
To enable the development of modern LLM-aware plagiarism detectors, we posit that the curation of high-quality, large-scale, type-varying plagiarism datasets is necessary. Although there exist a few plagiarism detection corpora (e.g., Potthast et al. (2013); Wahle et al. (2021a)), they are not comprehensive and up-to-date to adequately reflect the rapidly evolving writing patterns influenced by LLMs. In this work, therefore, we first develop the PlagBench corpus, which consists of 46.5K text pairs covering three types of LLM-aware plagiarism (i.e., verbatim, paraphrase, and summary cases) in three writing domains (i.e., abstract of scholarly paper, story, news article). In the context of PlagBench, we investigate two research questions: (RQ1) how well can LLMs generate paraphrase and summary plagiarism of given texts? and (RQ2) how well can LLMs detect three types of plagiarism? Figure 1 illustrates PlagBench creation processes, as well as our research questions. Specifically, we use three state-of-the-art (SOTA) chat-based LLMs, Llama2-70b-chat Touvron et al. (2023), GPT-3.5 Turbo [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo, and GPT-4 Turbo Achiam et al. (2023)) for PlagBench construction. Llama2-70b-chat, Llama3-70b-instruct, [^3] https://ai.meta.com/blog/meta-llama-3/ Mixtral-8x7B Jiang et al. (2024), GPT-3.5 Turbo, and GPT-4 Turbo are then used for the detection task. [^4] At the time of data generation, Llama3 models have not been released, and thus it was only employed in the detection process. Additionally, GPT-4o is excluded from both generation and detection tasks because it was unavailable during the time of the experiments.
Key findings of our experiments are as follows: (1) Among the three LLMs, GPT-3.5 Turbo produces the highest quality paraphrases and summaries, (2) LLMs like Llama3 and GPT-4, with just prompting, can outperform existing plagiarism checkers that are specifically trained for the task, and (3) LLMs generally have difficulty distinguishing summary plagiarism.
@@ 2.1 Plagiarism Detection Corpus
The Microsoft Research Paraphrase Corpus (MRPC) Dolan and Brockett (2005) and PAWS Zhang et al. (2019b) are among the most renowned datasets for paraphrase identification. Yet, they are not suitable for plagiarism detection tasks because it only covers one type. Additionally, their samples are annotated at the sentence level, which limits their applicability in real-world scenarios where plagiarism in general occurs at the paragraph level. More recently, Wahle et al. (2021b) constructed a paragraph long machine-paraphrased plagiarism dataset using three neural language models such as BERT, RoBERTa, and Longformer. Following, the authors replaced these three models with more recently released LLMs, GPT-3, and T5, and produce machine-paraphrased texts Wahle et al. (2022). There are many independent machine-summarized text pairs (e.g., Cohan et al. (2018); Narayan et al. (2018)) as well, but there exists no comprehensive dataset covering both paraphrases and summaries generated by modern LLMs.
@@ 2.2 Automatic Plagiarism Detection
There are two different formats of the automated plagiarism detection task: intrinsic and extrinsic plagiarism detection. While intrinsic plagiarism detection analyzes the document itself without resorting to the reference document, extrinsic plagiarism detection involves directly comparing the suspicious document against the reference to detect plagiarism instances Alzahrani et al. (2011). Intrinsic approaches use patterns such as linguistic differences within a document which indicates multiple authorship to detect plagiarism Gipp and Gipp (2014); Xiong et al. (2024); Potthast et al. (2010); Song et al. (2024); Stein et al. (2011). Due to high complexity of intrinsic plagiarism task, a majority of literature and software tools follow extrinsic plagiarism detection strategy. The extrinsic plagiarism detection task is generally divided into two subtasks: source retrieval (i.e., finding the most plausible pair of the source and the suspicious document) and plagiarism identification (i.e., classifying whether two documents are plagiarizing each other or not). Extrinsic approaches rely on similarity scores of two non-zero vectors transformed from the document pairs and apply certain thresholds Vani and Gupta (2017); Sağlam et al. (2024b); Potthast et al. (2010); Sağlam et al. (2024a); Lee et al. (2023a); Moravvej et al. (2023); Avetisyan et al. (2023). Our focus lies in the latter scope. We assume that we already have a text pair, the source document and the suspicious document obtained through PlagBench, and we concentrate on plagiarism identification.
3 Description of Automatic Plagiarism Detection
Suppose we have a source document $D_{src}$ and a suspicious document $D_{susp}$ that potentially plagiarizes $D_{src}$. Following the plagiarism detection task proposed by PAN, [^5] https://pan.webis.de we study three branches of plagiarism: verbatim, paraphrase, and summary. Verbatim plagiarism occurs when exact copies of words or phrases from $D_{src}$ are found in $D_{susp}$ without quotation marks. Paraphrase plagiarism, on the other hand, is when $D_{susp}$ rephrases $D_{src}$ using different words but retain the same meaning and structure without providing a citation. Synonymous substitution, word reordering and insertion/deletion are common paraphrasing techniques used by plagiarists Alvi et al. (2021). While both verbatim and paraphrase plagiarism cases tend to maintain the same length and the structure as the original text, summary plagiarism plagiarism involves succinctly summarizing $D_{src}$ to reuse its key points or ideas. As summaries leave out unnecessary information, they are prone to be significantly shorter than $D_{src}$. Unlike verbatim plagiarism, which can be effectively captured using simple string matching algorithms, paraphrase and summary plagiarism are more challenging to identify due to their limited lexical and syntactic similarities to the original text. Specifically, these two categories require careful attention to the meaning and context of the content rather than just high-level vocabulary overlaps.
4 PlagBench Creation
4.1 PlagBench Corpus Construction
Instruction tuning is a core training mechanism crucial for assisting LLMs in responding to human instructions. Once LLMs are pretrained with the objective of predicting subsequent words within large corpora, LLMs then undergo additional fine-tuning stage using diverse instruction datasets consisting of formatted instruction pairs. Instruction-tuned LLMs have demonstrated the ability to generalize to unseen instructions Raffel et al. (2020); Wei et al. (2021). Motivated by this, we leverage three prominent instruction-tuned LLMs, Llama2-70b-chat, GPT-3.5 Turbo, and GPT-4 Turbo, to create the PlagBench corpus. Our experiment design is as follows: given one source document $D_{src}$, three independent LLMs are prompted to rewrite it according to our description. We then consider their outputs as $D_{susp}$. Specifics of source document collection and suspicious document generation are described below.
4.1.1 Source Document Collection
@@
Although plagiarism can occur in any type of writing, we target specific areas such as scholarly or creative works, where originality and intellectual property are highly valued. Li et al. (2023) present a collection of human-authored and machine-authored corpora from diverse writing tasks. Among their selections of corpora, we carefully choose three public English datasets: (1) SciXGen Chen et al. (2021) consisting of 200K+ abstracts of scientific articles (for scientific writing), (2) ROCStories Mostafazadeh et al. (2016) consisting of 50k five-sentence commonsense stories (for story writing); and (3) TLDR [^6] https://huggingface.co/datasets/JulesBelveze/tldr_news consisting of 7K+ TLDR tech newsletters (for news article writing). Their samples contain one human-written text and multiple machine-generated variations, all sharing either the same topic or text source as the human-written piece. For the purpose of the study, we only use the human-written text containing at least 5 sentences as the source text. [^7] According to our pilot experiment, paraphrases and summaries generated based on documents shorter than 5 sentences tend to too similar to distinguish even from human lens. We randomly select 2,400 samples from SciXGen and ROCStories corpora and 1,300 from TLDR for suspicious document generation. The sample size of TLDR is smaller than the other two datasets because content of TLDR tends to be short.
4.1.2 Suspicious Document Generation
@@
Instruction-tuned LLMs have demonstrated strong abilities on text rephrasing Lingard (2023); Zhang et al. (2024); Tang et al. (2023). Hence, we employ Llama2-70b-chat, GPT-3.5 Turbo, and GPT-4 Turbo models to generate texts corresponding to two plagiarism labels (i.e., paraphrase and summary plagiarism) respectively. We do not use these models for verbatim plagiarism, as generating suspicious documents can be easily accomplished by simple copy-and-paste without any modification. Table 4 shows our hand-crafted prompt templates, each tailored to transform a single source document according to the provided task descriptions. As opposed to our expectations, our pilot experiment hints that some models, especially the Llama models, have difficulties in providing correct paraphrases or summaries on the fly. Thus, following findings that in-context learning is very effective in enabling LLMs to perform a new task without additional training Liu et al. (2021); Min et al. (2022), we provide three demonstrations, one from each writing domains, inside the prompt.
@@
In plagiarism-free scenarios, existing datasets (e.g., Potthast et al. (2013), Foltỳnek et al. (2020)) paired random or dissimilar documents. However, this often results in pairs that are clearly different and thus easily identifiable. More challenging pairs are likely to occur when two documents share the same key topics and domain type, but their details do not overlap. Consequently, we instruct LLMs to write continuations based on provided keywords, domain information, and the first two sentences of the source document. Two meaningful keywords are automatically extracted using KeyBERT, [^8] https://maartengr.github.io/KeyBERT/index.html a package leveraging contextual embeddings from BERT Devlin et al. (2019) to generate keyphrases resembling the provided document. We do not perform in-context learning for plagiarism-free case generation because LLMs, originally trained for next token prediction, handled the task well without needing demonstrations.

### v2

@@ Abstract
Recent studies have raised concerns about the potential threats large language models (LLMs) pose to academic integrity and copyright protection. Yet, their investigation is predominantly focused on literal copies of original texts. Also, how LLMs can facilitate the detection of LLM-generated plagiarism remains largely unexplored. To address these gaps, we introduce PlagBench, a dataset of 46.5K synthetic text pairs that represent three major types of plagiarism: verbatim copying, paraphrasing, and summarization. These samples are generated by three advanced LLMs. We rigorously validate the quality of PlagBench through a combination of fine-grained automatic evaluation and human annotation. We then utilize this dataset for two purposes: (1) to examine LLMs’ ability to transform original content into accurate paraphrases and summaries, and (2) to evaluate the plagiarism detection performance of five modern LLMs alongside three specialized plagiarism checkers. Our results show that GPT-3.5 Turbo can produce high-quality paraphrases and summaries without significantly increasing text complexity compared to GPT-4 Turbo. However, in terms of detection, GPT-4 outperforms other LLMs and commercial detection tools by 20%, highlights the evolving capabilities of LLMs not only in content generation but also in plagiarism detection. Data and source code are available at https://github.com/Brit7777/plagbench.
@@ 1 Introduction
Plagiarism occurs when someone uses another person’s work, ideas, or expressions without proper acknowledgment. It is one of the most frequently scrutinized issues in writing tasks, which leads to a violation of intellectual property rights and academic integrity. Traditionally, plagiarism research has focused on human-written text. However, the rise of large language models (LLMs) has transformed the landscape of plagiarism Pudasaini et al. (2024). For instance, users can exploit these models to obfuscate original content through paraphrases or summaries, effectively bypassing traditional plagiarism detection systems Krishna et al. (2024); Sadasivan et al. (2023); Weber-Wulff et al. (2023); Elkhatat et al. (2021). Moreover, many recent studies Carlini et al. (2021); Tirumala et al. (2022); Zhang et al. (2023); Carlini et al. (2023) have shown that LLMs can memorize portions of their training data and reproduce them during text generation. These models are not limited to producing verbatim copies; they can also generate paraphrased or elongated versions of unique content Lee et al. (2023b).
@@
The increasing prevalence of plagiarism in LLMs raises two critical research questions: (RQ1) how well can LLMs generate paraphrase and summary plagiarism of given texts? and (RQ2) how well can LLMs detect three types of plagiarism–i.e., verbatim, paraphrase, and summary plagiarisms?. While prior research has explored LLMs’ applications to automated paraphrase Witteveen and Andrews (2019); Wahle et al. (2022) and summary Lingard (2023); Goyal et al. (2022); Liu et al. (2023) generation, these studies are often limited in scope because most of them focus on individual domains or specific models. A holistic and fine-grained analysis of diverse domains and models is crucial to developing robust mechanisms for detecting and mitigating plagiarism in AI-generated content. Additionally, LLMs have demonstrated impressive performance in various natural language processing (NLP) tasks Chang et al. (2024), motivating us to investigate their potential in automatic plagiarism detection. Current plagiarism detectors are primarily optimized to catch literal copies or superficial paraphrasing by humans and may not be sufficient to detect nuanced rewritings of LLMs. Evaluating LLMs’ effectiveness in plagiarism detection can offer critical insights into the limitations of current systems and help guide the development of more advanced, robust solutions.
To answer these questions, we present PlagBench, a large-scale corpus containing 46.5K text pairs with three plagiarism categories: verbatim, paraphrase, and summary plagiarism. These samples cover three writing domains (i.e., abstract of scholarly paper, story, news article) and are synthetically generated by three advanced, closed-source and open-source LLMs (Llama2-70b-chat Touvron et al. (2023), GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo, and GPT-4 Turbo Achiam et al. (2023)). To acquire the gold standard for automatic plagiarism detection tasks, we employ a comprehensive integration of both automatic and manual evaluation. This final subset is then used for evaluation of five popular LLMs (Llama2-70b-chat, Llama3-70b-instruct [^2] https://ai.meta.com/blog/meta-llama-3/, Mixtral-8x7B Jiang et al. (2024), GPT-3.5 Turbo, and GPT-4 Turbo) and three specialized plagiarism detectors (GPTZero [^3] https://gptzero.me/plagiarism-checker, Prepostseo [^4] https://www.prepostseo.com/plagiarism-checker, and a detector proposed by Lee et al. (2023b)), two of which are commercial-ready. Figure 1 illustrates a visual representation of our dataset construction processes and the proposed RQs.
Key contributions of this work are as follows: (1) PlagBench is the first dataset that covers a wide range of plagiarism examples generated by state-of-the-art LLMs; (2) we showcase the strengths and weaknesses of LLM-generated plagiarism cases across domains; (3) our results show the potential of LLMs facilitating plagiarism detection, although they generally have difficulties in distinguishing the type of summary plagiarism.
@@ 2.1 Plagiarism Detection Corpus
The Microsoft Research Paraphrase Corpus (MRPC) Dolan and Brockett (2005) and PAWS Zhang et al. (2019b) are among the most renowned datasets for paraphrase identification. Yet, they are not suitable for plagiarism detection tasks because they only cover one type. Additionally, their samples are annotated at the sentence level, which limits their applicability in real-world scenarios where plagiarism in general occurs at the paragraph level. More recently, Wahle et al. (2021) constructed a paragraph-long machine-paraphrased plagiarism dataset using three neural language models such as BERT Devlin et al. (2019a), RoBERTa Liu (2019), and Longformer Beltagy et al. (2020). Following, the authors replaced these three models with more recently released LLMs, GPT-3, and T5, and produced machine-paraphrased texts Wahle et al. (2022). There are many independent machine-summarized text pairs (e.g., Cohan et al. (2018); Narayan et al. (2018)) as well, but there exists no comprehensive dataset covering both paraphrases and summaries generated by modern LLMs.
@@ 2.2 Automatic Plagiarism Detection
There are two formats of the automated plagiarism detection task: intrinsic and extrinsic plagiarism detection. While intrinsic plagiarism detection analyzes the document itself without resorting to the reference document, extrinsic plagiarism detection involves directly comparing the suspicious document against the reference to detect plagiarism instances Alzahrani et al. (2011). Intrinsic approaches use patterns such as linguistic differences within a document which indicates multiple authorship to detect plagiarism Gipp and Gipp (2014); Xiong et al. (2024); Potthast et al. (2010); Song et al. (2024); Stein et al. (2011). Due to the high complexity of intrinsic plagiarism tasks, many popular detection tools follow extrinsic plagiarism detection strategies. The extrinsic plagiarism detection task is generally divided into two subtasks: source retrieval (i.e., finding the most plausible pair of the source and the suspicious document) and plagiarism identification (i.e., classifying whether two documents are plagiarizing each other or not). These approaches rely on similarity scores of two non-zero vectors transformed from the document pairs and apply certain thresholds Vani and Gupta (2017); Sağlam et al. (2024b); Potthast et al. (2010); Sağlam et al. (2024a); Lee et al. (2023a); Moravvej et al. (2023); Avetisyan et al. (2023). Our focus lies in the latter scope. We assume we already have the text pair, the source document, and the suspicious document obtained through PlagBench, and we concentrate on plagiarism identification.
3 PlagBench Corpus Generation
3.1 Corpus Construction
@@
Source Document Collection. Although plagiarism can occur in any type of writing, we target specific areas such as scholarly or creative works, where originality and intellectual property are highly valued. Li et al. (2023) present a collection of human-authored and machine-authored corpora from diverse writing tasks. Among their selections of corpora, we carefully choose three public English datasets: (1) SciXGen Chen et al. (2021) consisting of 200K+ abstracts of scientific articles (for scientific writing), (2) ROCStories Mostafazadeh et al. (2016) consisting of 50k five-sentence commonsense stories (for story writing); and (3) TLDR [^5] https://huggingface.co/datasets/JulesBelveze/tldr_news consisting of 7K+ TLDR tech newsletters (for news article writing). Their samples contain one human-written text and multiple machine-generated variations, all sharing either the same topic or text source as the human-written piece. For the purpose of the study, we only use the human-written text containing at least 5 sentences as the source text. [^6] According to our pilot experiment, paraphrases and summaries generated based on documents shorter than 5 sentences tend to be too similar to distinguish even from the human lens. We randomly select 2,400 samples from SciXGen and ROCStories corpora and 1,300 from TLDR for suspicious document generation. The sample size of TLDR is smaller than the other two datasets because the content of TLDR tends to be short.
@@
Suspicious Document Generation. Instruction-tuned LLMs have demonstrated strong abilities in text generation that are aligned with users’ natural language commands Lingard (2023); Zhang et al. (2024); Tang et al. (2023). Hence, we employ Llama2-70b-chat, GPT-3.5 Turbo, and GPT-4 Turbo models to generate texts corresponding to two plagiarism labels (i.e., paraphrase and summary plagiarism), respectively. We do not use these models for verbatim plagiarism, as generating suspicious documents can be easily accomplished by simple copy-and-paste without any modification. Table 8 shows our hand-crafted prompt templates, each tailored to transform a single source document according to the provided task descriptions.
@@
In plagiarism-free scenarios, existing datasets (e.g., Potthast et al. (2013), Foltỳnek et al. (2020)) pair random or dissimilar documents. However, this often results in pairs that are clearly different and thus easily identifiable. More challenging pairs are likely to occur when two documents share the same key topics and domain type, but their details do not overlap. Consequently, we instruct LLMs to write continuations based on provided keywords, domain information, and the first two sentences of the source document. Two meaningful keywords are automatically extracted using KeyBERT, [^7] https://maartengr.github.io/KeyBERT/index.html a package leveraging contextual embeddings from BERT Devlin et al. (2019b).